What is the limiting reagent when 0.50 mol of Cr and 1.0 mol of H3PO4 react according
to the following chemical equation?
2Cr + 2H3PO4 → 2CrPO4 + 3H2
If 0.20 mol of CrPO4 is recovered from the reaction described above, what is the percent yield?
1
Expert's answer
2019-06-18T05:46:04-0400

Limiting reagent is reagent that is in a smaller molar amount in a chemical reaction. Limiting reagent limits the amount of products that can be made. In the reaction shown above the limiting reagent is Cr (0.50 mole < 1.0 mole).

"percent \space yield = \frac {actual \space yield }{theoretical \space yield} \cdot 100 \%"

"actual \space yield = 0.20 \space mol \cdot \frac {146.97 g/mol} {1mol} = 29 g"


Calculate theoretical yield:


"mol \space of \space CrPO_4 = 0.50 \space mol \space Cr
 \cdot \frac {2mol \space CrPO_4}{2 mol \space Cr}=0.50 mol"


"theoretical \space yield = mass \space of \space CrPO_4 = 0.50 \cdot 146.97 \space g/mol = 73 \space g"


"percent \space yield = \frac {29 g}{73g} \cdot 100 \% =40 \%"